# PixWizard: Versatile Image-to-Image Visual Assistant with Open-Language Instructions

### v2

@@ Abstract
This paper presents a versatile image-to-image visual assistant, PixWizard, designed for image generation, manipulation, and translation based on free-from language instructions. To this end, we tackle a variety of vision tasks into a unified image-text-to-image generation framework and curate an Omni Pixel-to-Pixel Instruction-Tuning Dataset. By constructing detailed instruction templates in natural language, we comprehensively include a large set of diverse vision tasks such as text-to-image generation, image restoration, image grounding, dense image prediction, image editing, controllable generation, inpainting/outpainting, and more. Furthermore, we adopt Diffusion Transformers (DiT) as our foundation model and extend its capabilities with a flexible any resolution mechanism, enabling the model to dynamically process images based on the aspect ratio of the input, closely aligning with human perceptual processes. The model also incorporates structure-aware and semantic-aware guidance to facilitate effective fusion of information from the input image. Our experiments demonstrate that PixWizard not only shows impressive generative and understanding abilities for images with diverse resolutions but also exhibits promising generalization capabilities with unseen tasks and human instructions. The code and related resources are available at https://github.com/AFeng-x/PixWizard.
@@ 1 Introduction
Figure 1: Task Overview of our Omni Pixel-to-Pixel Instruction-tuning Dataset for PixWizard.
@@
$\bullet$ 3. Architecture Design: For a robust visual assistant, the architecture and scalability of the foundational model are crucial. In this work, we use the flow-based Diffusion Transformer (DiT) (Ma et al., 2024) as our base model. Compared to traditional diffusion models, flow-based models offer greater versatility and stability in modeling data distributions by learning conditional velocity fields. The DiT architecture further enhances scalability and is well-suited for incorporating conditional information. Building on this foundation, we introduce several unique and effective components. Specifically, we extend the dynamic partitioning and padding scheme to enable the model to handle input images of any resolution, closely aligning with human perceptual processes. Additionally, we introduce structure-aware and semantic-aware guidance, allowing the model to effectively follow multimodal instructions (images and open-ended user prompts) for a variety of manipulations.
Experimental results show that PixWizard delivers competitive performance compared to task-specific vision models and surpasses current state-of-the-art general visual models. More importantly, our model handles tasks and instruction prompts it has not encountered during training, demonstrating impressive generalization capabilities. This further highlights PixWizard’s strength as a powerful interactive image-to-image visual assistant.
@@ 2 Omni Pixel-to-Pixel Instruction-Tuning Dataset
To equip our image-to-image visual assistant with comprehensive capabilities in image generation, manipulation, and translation, we compiled a multi-task training dataset for visual instruction tuning, consisting of 30 million instances across seven primary domains, as illustrated in Fig. 1. To our knowledge, this is the largest and user-friendly image-instruction-image triplet dataset, built from both open-source and in-house data, filtered with the help of MLLMs and manual review.
Image Restoration. We incorporate low-level data to restore images degraded by various environmental or technical factors. This section utilizes a wide range of open-source datasets covering key restoration tasks, including (1) Denoising, (2) Deraining, (3) Demoireing, (4) Dehazing, (5) Deblurring, (6) Desnowing, (7) Deshadowing, (8) Low-Light Enhancement, (9) Face Restoration, (10) Watermark Removal, and (11) Super Resolution. Since both the inputs and outputs are inherently defined in the RGB space, these tasks can be seamlessly unified by our PixWizard model without extra any transformations. All open-source datasets we use are provided in Sec. B.1.
@@
Inpainting involves filling in missing parts of an image with new or modified content. To create inpainting instances, we apply random black or white masks to different regions of the original images. These masks come in various shapes, including circles, rectangles, and free-form patterns, and are randomly placed on the images, resulting in a wide range of occluded areas for inpainting.
Outpainting (Image extrapolation) extends an image’s content beyond its boundaries. To create outpainting instances, we randomly crop the central part of the image, while masking the surrounding areas with black or white. The cropping is not limited to square shapes but includes rectangles of varying proportions. This approach ensures that the outpainting task covers a range of extension scenarios, challenging the model to generate coherent content beyond the original image boundaries.
Text-to-Image Generation. Image-to-image and text-to-image are two distinct tasks. The former requires an additional input image as a condition, while the latter does not. Existing instruction-tuning methods transform pretrained text-to-image models into image-to-image systems, but they fail to retain their original text-to-image capabilities. To unify both tasks within a single framework, we propose a "drawing" strategy to preserve the text-to-image capabilities. Specifically, we introduce an additional input alongside the language instructions: an image that is either entirely white or black. This setup simulates a blank canvas, allowing the model to "draw" the generated images based on the provided text prompts. This approach distinguishes our model from previous text-to-image systems. All the image-text pairs we use are sourced from our own collection of high-quality images from the Internet, with captions generated using MLLMs assistance.
@@ 3 PixWizard
PixWizard is a versatile image-to-image model fine-tuned on our custom-built omni pixel-to-pixel instruction-tuning dataset. In this section, we present the details of the PixWizard framework from the perspectives of model architecture (as shown in Fig. 2) and training strategies.
@@ 3.2 Architecture
Text Encoders. We begin by using Gemma-2B (Team et al., 2024) as the text embedder in PixWizard to encode text prompts. However, in multi-task learning, relying solely on text instructions is insufficient to guide the model in accurately executing user commands. To better guide the generation process for the correct task, we incorporate the CLIP text encoder (Radford et al., 2021). Global average pooling is applied to the CLIP text embeddings to obtain a coarse-grained text representation, which is passed through an MLP-based task embedder to generate the task embedding. This embedding is then integrated into the PixWizard Block by adding it to the timestep embeddings through a modulation mechanism. As shown in Fig. 3, this approach adaptively clusters similar task instructions in the latent space while separating those from different tasks, which helps guide the model’s generation process in the correct task direction.
Figure 3: t-SNE visualization of the global text embeddings. Each dot represents a human instruction.
Structural-Aware Guidance. To effectively capture the overall structural features of the input image condition, we begin by encoding the image using a variational autoencoder (VAE) (Kingma & Welling, 2013) from SDXL (Podell et al., 2023). Next, we concatenate the image latent with the noise latent along the channel dimension. Following (Brooks et al., 2023), additional input channels are added to the patch embedder, with the weights for these new channels initially set to zero.
@@
However, inputting all image tokens into the attention layer often leads to significant computational demands, and we also found that not all semantic tokens are relevant to the specific task. To address this, we introduce the Task-Aware Dynamic Sampler, which is designed to select the most relevant semantic tokens for each task. This sampler uses a lightweight ranking network consisting of four linear layers and activation functions. Inspired by DynamicViT (Rao et al., 2021), we employ a technique that maps image tokens to both local and global features. Furthermore, we integrate task embeddings ${x_{task}}$ to help the sampler identify the tokens that are most relevant to the task. The computational process is formulated as:
@@
where the output of GumbelSoftmax is a multi-hot tensor that represents the mask of the retained tokens. $\odot$ is the Hadamard product, indicating that the top $K$ tokens by importance score are weighted by 1 and thus retained, while the remaining $(N-K)$ tokens are weighted by zero and discarded. Finally, we equip each layer of the transformer block with an independent task-aware dynamic sampler. This approach not only helps capture the most relevant semantic features needed by each layer to fulfill the requirements of different tasks but also reduces the computational cost in the attention process.
3.3 Any Resolution
PixWizard inherits the dynamic partitioning and padding scheme introduced by (Zhuo et al., 2024), allowing the model to handle images of any resolution and aspect ratio during both fine-tuning and inference. However, in practice, the required resolutions for different tasks can vary significantly. To support more flexible handling of arbitrary resolutions while preserving the original resolution as much as possible, we use $[512^{2},768^{2},1024^{2}]$ as resolution centers to generate candidate patch partitions. During training, we group data with similar resolutions into buckets, ensuring that sequence lengths within each batch are comparable, minimizing padding tokens, and improving training efficiency. During inference, by incorporating NTK-Aware Scaled RoPE (Peng & Quesnelle, 2023) and sandwich normalization, PixWizard demonstrates exceptional resolution extrapolation.
@@ 3.4 Two-Stage Training and Data Balancing Strategies
To unlock the model’s potential and improve its performance on tasks with smaller datasets, we propose a two-stage training and data-balancing strategy. S1: At this stage, we initialize the model’s weights by combining the weights of a pretrained text-to-image model with randomly initialized weights for the newly added modules. We begin by selecting tasks with smaller datasets and assigning each dataset a sampling weight to increase its data volume. This weight determines how many times the dataset is repeated during a single epoch. Using this method, each task has approximately 20k data points. We then randomly select training samples from other tasks to match this scale, creating our first-stage training dataset. The training process lasts for a total of 4 epochs. S2: In the second stage, we initialize the model using the weights obtained from the first stage and combine all collected data for further training. To balance tasks, we manually assign sampling weights to each dataset, randomly selecting data if a weight is less than 1.0. We also include text-to-image data at a 1:1 ratio with other tasks, resulting in our second-stage training dataset. At this stage, the total amount of training data reaches up to 20 million samples.
@@ 4.1 First Section Results
PixWizard 0.287 0.288 36.76 19.65 38.67 0.957 31.43 0.917 46.44 36.49
Settings. For image restoration, we follow previous works (Conde et al., 2024; Potlapalli et al., 2024) and prepare datasets for various restoration tasks during training. For evaluation, we first select two representative benchmarks: Rain100L (2017) for deraining and SIDD (2018) for denoising. Additionally, to further assess performance on other restoration tasks and evaluate zero-shot capabilities, more details and results are provided in the Sec. D.1.
@@
Dense image prediction tasks are evaluated on the three vision tasks: ADE20k (2017b) for semantic segmentation, NYUv2 (2012) and SUNRGB-D (2014) for monocular depth estimation, and NYU-Depth v2 (2012) for surface normal estimation. For semantic segmentation, we assign labels by identifying the nearest neighbor RGB color value, and accuracy is evaluated using the Mean Intersection over Union (mIoU) metric. For monocular depth estimation, we average the output image across the three channels and apply the inverse of the linear transformation used during training to obtain depth estimates within the range of $[0,10]$ meters. Accuracy is evaluated using the Root Mean Square Error (RMSE). For surface normal estimation, we recover the corresponding normal vectors from the output image and use the Mean Angle Error to assess accuracy.
Results. Table 1 presents a comprehensive performance comparison with recent state-of-the-art (SOTA) task-specific and all-in-one methods. As shown in the results, despite denoising and deraining data making up only a small portion of the overall training set, our method outperforms other unified methods and even surpasses some task-specific approaches. In the image grounding task, PixWizard significantly outperforms the diffusion-based generalist model InstructDiffusion by 4.8 cIoU on RefCOCO (val). However, there is still room for improvement compared to other highly specialized models. Furthermore, as shown in Fig.6, PixWizard supports flexible instructions, allowing it to not only highlight and visualize the target object directly on the image but also generate the corresponding binary mask. This highlights its strong performance in real-world interactions and practical applications. Additional quantitative evaluation results are provided in the Sec. D.2.
@@
For dense prediction tasks, PixWizard performs competitively against both generalist and task-specific baselines across all three tasks. In depth estimation on the NYUv2 test set, PixWizard achieves a 10.0% improvement in RMSE compared to Unified-IO and performs similarly to Painter and InstructCV. For semantic segmentation, PixWizard surpasses Unified-IO by +11.05 in mIoU, though it still trails behind other methods. Additionally, Fig. 7 provides examples of PixWizard’s outputs. As shown, by supplying the corresponding task-specific prompt for the same image, we can easily generate the respective condition visualization, underscoring PixWizard’s significant practical value.
@@ 4.2 Second Section Results (Image Editing)
Settings. We evaluate PixWizard across two benchmarks, the MagicBrush Test (Zhang et al., 2024a) and the Emu Edit Test (Sheynin et al., 2024), to assess its effectiveness in image editing capabilities. For a fair comparison, we primarily compare it with instruction-guided image editing methods, including InstructPix2Pix, MagicBrush, Emu Edit, and UltraEdit. Consistent with Emu Edit, we use L1 distance, CLIP image similarity, DINO similarity, CLIP text-image similarity, and CLIP text-image direction similarity as metrics.
@@
Results. Table 2 presents our results compared to the baselines. The findings show that our model consistently outperforms InstructPix2Pix, MagicBrush, and UltraEdit in automatic metrics and achieves comparable performance to state-of-the-art method Emu Edit. Qualitative comparisons are provided in Fig. 8. Our model precisely identifies the editing region while preserving the rest of the pixels, and it demonstrates the best understanding of the given instructions.
@@ 4.3 Third Section Results (Image Generation)
In this section, we focus on evaluating the effectiveness of PixWizard’s generation capabilities. Specifically, we assess its performance across four key tasks: classic text-to-image generation, controllable image generation, inpainting, and outpainting.
Settings. For controllable image generation, we evaluated PixWizard’s ability to generate images based on two specific conditions: canny edge maps and depth maps. Following ControlNet++ (Li et al., 2024), we assessed controllability by measuring the similarity between the input conditions and the corresponding features extracted from the generated images. Specifically, we used RMSE for depth map control and F1-Score for canny edges. Additionally, to evaluate the quality of the generated images and their alignment with the input text, we reported FID (Fréchet Inception Distance) and CLIP-Score metrics. All experiments were conducted at a resolution of $512\times 512$.
For image inpainting, we use latent diffusion settings (Rombach et al., 2022b) to measure FID and LPIPS, which assess the quality of the generated samples when 40-50% of the image area needs to be inpainted. For image extrapolation (outpainting), we follow MaskGIT (Chang et al., 2022a) settings, extending the image by 50% to the right and comparing performance against common baselines using FID and Inception Score (IS). Both tasks are evaluated on 30,000 image crops, each sized $512\times 512$, from the Places dataset (Zhou et al., 2017a).
For the text-to-image generation task, we use two primary evaluation methods. First, we visually present examples of images generated by PixWizard. Additionally, we calculate two automatic evaluation metrics: the Human Preference Score (HPS) v2 (Wu et al., 2023) and the standard Zero-shot FID-30K on the MS-COCO dataset (Lin et al., 2014).
@@
Controllable Generation Results. Without the need for separate training for each model, PixWizard is an all-in-one solution capable of handling multiple conditions. As shown in Table 3, PixWizard achieves the highest controllability and best image quality under depth conditions, while also being comparable to current separate models in image-text alignment. Fig. 9 presents several visual samples, demonstrating the effectiveness of our approach.
Figure 10: Visualization results of Inpainting and Outpainting.
Inpainting Results. The comparison with other inpainting approaches in Table 3 shows that PixWizard improves overall image quality, as indicated by FID and LPIPS. The qualitative examples in Fig. 10 further demonstrate PixWizard’s effectiveness in generating coherent content. We attribute this to PixWizard’s ability to "paint" on a blank canvas, which greatly enhances its capacity to identify masked areas and generate coherent content within them.
Building on this strong inpainting capabilities,, PixWizard allows for more precise image editing tasks: (i) Remove Anything tackles the object removal problem (Criminisi et al., 2003; 2004; Elharrouss et al., 2020), enabling users to remove specific objects while keeping the result visually seamless. The process consists of two steps: identifying and removing. As shown in Fig. 11, the user first instructs PixWizard on which object to locate, and PixWizard generates a target mask. The mask is then applied to the original image, and PixWizard fills the area with appropriate background details. (ii) Replace Anything lets users swap out any object in an image. The process is similar to Remove Anything, but instead of just removing the object, the model replaces it with the specified object while ensuring the background remains visually consistent. (iii) Add Anything enables users to insert any object into an image and place it wherever they like. The user adds a mask to the desired area and provides a text prompt specifying what to add. PixWizard then leverages its powerful inpainting ability to generate the requested content.
Figure 11: Visualization results of Remove, Replace and Add Anything.
@@
Text-to-Image Results. As shown in the quantitative comparison results in Table 3, PixWizard achieves a FID score of 9.56 when tested for zero-shot performance on the COCO dataset. Although some models achieve a lower FID, they focus solely on text-to-image tasks and rely on significantly more training resources. Additionally, we evaluated the Human Preference Score (HPS v2), a robust benchmark for assessing human preferences in text-to-image synthesis. PixWizard performed well, delivering image quality comparable to popular text-to-image generators. We also provide visual samples in Fig. 12. PixWizard supports high-resolution image synthesis, up to $1024\times 1024$, with any resolution and aspect ratio.
Figure 12: Text-to-image samples generated by PixWizard.

### v3

@@ Abstract
This paper presents a versatile image-to-image visual assistant, PixWizard, designed for image generation, manipulation, and translation based on free-from language instructions. To this end, we tackle a variety of vision tasks into a unified image-text-to-image generation framework and curate an Omni Pixel-to-Pixel Instruction-Tuning Dataset. By constructing detailed instruction templates in natural language, we comprehensively include a large set of diverse vision tasks such as text-to-image generation, image restoration, image grounding, dense image prediction, image editing, controllable generation, inpainting/outpainting, and more. Furthermore, we adopt Diffusion Transformers (DiT) as our foundation model and extend its capabilities with a flexible any resolution mechanism, enabling the model to dynamically process images based on the aspect ratio of the input, closely aligning with human perceptual processes. The model also incorporates structure-aware and semantic-aware guidance to facilitate effective fusion of information from the input image. Our experiments demonstrate that PixWizard not only shows impressive generative and understanding abilities for images with diverse resolutions but also exhibits generalization capabilities with unseen tasks and human instructions. The code and related resources are available at https://github.com/AFeng-x/PixWizard.
@@ 1 Introduction
Figure 1: Task Overview of the Omni Pixel-to-Pixel Instruction-tuning Dataset for PixWizard.
@@
$\bullet$ 3. Architecture Design: For a robust visual assistant, the architecture and scalability of the foundation model are crucial. In this work, we use the flow-based Diffusion Transformer (DiT) (Ma et al., 2024), which offer great versatility and stability in modeling data distributions by learning conditional velocity fields. The DiT architecture further enhances scalability and is well-suited for incorporating conditional information. Building on this foundation, we introduce several key components. We extend the dynamic partitioning and padding scheme (Zhuo et al., 2024) to handle input images of any resolution, aligning closely with human perception. Additionally, we implement structure-aware and semantic-aware guidance, enabling the model to follow multimodal instructions (images and user prompts) for a wide range of manipulations.
Experimental results show that PixWizard achieves competitive performance on certain tasks compared to task-specific vision models, and outperforms current state-of-the-art general visual models in most tasks, delivering its strong overall multi-task performance. More importantly, our model handles tasks and instruction prompts it has not encountered during training, demonstrating promising generalization capabilities. This further highlights PixWizard’s strength as a powerful interactive image-to-image visual assistant.
@@ 2 Omni Pixel-to-Pixel Instruction-Tuning Dataset
To equip our image-to-image visual assistant with comprehensive capabilities in image generation, manipulation, and translation, we compiled a multi-task training dataset for visual instruction tuning, consisting of 30 million instances across seven primary domains, as illustrated in Fig. 1. This is the user-friendly image-instruction-image triplet dataset, built from both open-source and in-house data, filtered with the help of MLLMs and manual review.
Image Restoration. We incorporate low-level data to restore images degraded by various environmental or technical factors. This section utilizes a wide range of open-source datasets covering key restoration tasks, including (1) Denoising, (2) Deraining, (3) Demoireing, (4) Dehazing, (5) Deblurring, (6) Desnowing, (7) Deshadowing, (8) Low-Light Enhancement, (9) Face Restoration, (10) Watermark Removal, and (11) Super Resolution. Since both the inputs and outputs are inherently defined in the RGB space, these tasks can be seamlessly unified by our PixWizard model without any extra transformations. All open-source datasets we use are provided in Sec. B.1.
@@
Inpainting involves filling in missing parts of an image with new or modified content. To create inpainting instances, we apply random black or white masks to different regions of the original images. These masks come in various shapes, including circles, rectangles, and free-form patterns, and are randomly placed on the images, resulting in a wide range of occluded areas for inpainting. Outpainting (Image extrapolation) extends an image’s content beyond its boundaries. To create outpainting instances, we randomly crop the central part of the image, while masking the surrounding areas with black or white. The cropping is not limited to square shapes but includes rectangles of varying proportions. This approach ensures that the outpainting task covers a range of extension scenarios, challenging the model to generate coherent content beyond the original image boundaries.
Text-to-Image Generation. Image-to-image and text-to-image are distinct tasks: the former requires an additional input image as a condition, while the latter does not. Existing instruction-tuning methods adapt pretrained text-to-image models for image-to-image tasks, but often lose the original text-to-image capabilities. To unify both tasks within a single framework, we propose a "drawing" strategy that preserves text-to-image functionality. Specifically, we introduce an additional input—a fully white or black image—alongside language instructions. This simulates a blank canvas, allowing the model to "draw" images based on the text prompts. This approach differentiates our model from previous text-to-image systems.
@@ 3 PixWizard
In this section, we present the details of the PixWizard framework from the perspectives of model architecture (as shown in Fig. 2) and training strategies.
@@ 3.2 Architecture
Text Encoders. We begin by using Gemma-2B (Team et al., 2024) as the text embedder in PixWizard to encode text prompts. However, in multi-task learning, relying solely on text instructions is insufficient to guide the model in accurately executing user commands. To better guide the generation process for the correct task, we incorporate the CLIP text encoder (Radford et al., 2021). Global average pooling is applied to the CLIP text embeddings to obtain a coarse-grained text representation, which is passed through an MLP-based task embedder to generate the task embedding. This embedding is then integrated into the PixWizard Block by adding it to the timestep embeddings through a modulation mechanism. In Sec. C.2, we show the t-SNE visualization to further illustrate the effectiveness of the task embedding.
Structural-Aware Guidance. To effectively capture the overall structural features of the input image condition, we begin by encoding the image using a variational autoencoder (VAE) (Kingma & Welling, 2013) from SDXL (Podell et al., 2023). Next, we concatenate the image latent with the noise latent along the channel dimension (Saharia et al., 2022c; a). Following (Brooks et al., 2023), additional input channels are added to the patch embedder, with the weights for these new channels initially set to zero.
@@
However, inputting all image tokens into the attention layer can lead to significant computational demands, and not all semantic tokens are relevant to the specific task. To address this, we introduce the Task-Aware Dynamic Sampler, designed to select the most relevant semantic tokens for each task. This sampler uses a lightweight ranking network with four linear layers and activation functions. Inspired by DynamicViT (Rao et al., 2021), we map image tokens to both local and global features. Additionally, task embeddings ($x_{task}$) are integrated to help the sampler identify the tokens most relevant to the task. The computational process is formulated as:
@@
where the output of GumbelSoftmax is a multi-hot tensor representing the mask of the retained tokens. $\odot$ denotes the Hadamard product, where the top $K$ tokens by importance score are weighted by 1 and retained, while the remaining $(N-K)$ tokens are weighted by zero and discarded. Each layer is equipped with an independent task-aware dynamic sampler. This approach not only captures the most relevant semantic features needed by each layer for different tasks but also reduces the computational cost in the attention process.
@@ 3.3 Two-Stage Training and Data Balancing Strategies
To unlock the model’s potential and improve its performance on tasks with smaller datasets, we propose a two-stage training and data-balancing strategy. S1: In the first stage, we initialize the model by combining the weights of a pre-trained Lumina-Next-T2I (Zhuo et al., 2024) with randomly initialized weights for the newly added modules. We prioritize tasks with smaller datasets, assigning each a sampling weight to increase its data volume. This weight determines how many times the dataset is repeated during an epoch. Using this method, each task achieves approximately 20k data points. We then randomly sample from other tasks to match this scale, creating our first-stage training dataset, with training spanning 4 epochs. S2: In the second stage, we initialize the model using the weights from the first stage and combine all the collected data for further training. To balance tasks, we assign manual sampling weights to each dataset, randomly selecting data when a weight is less than 1.0. We also include text-to-image data at a 1:1 ratio with other tasks, resulting in a second-stage training dataset. At this stage, the total dataset reaches up to 20 million samples.
@@ 4.1 First Section Results
PixWizard 0.287 0.291 32.76 19.65 38.67 0.957 31.43 0.917 46.44 36.49
Settings. For image restoration, we follow previous works (Conde et al., 2024; Potlapalli et al., 2024) and prepare datasets for various restoration tasks during training. For evaluation, we first select two representative benchmarks: Rain100L (2017) for deraining and SIDD (2018) for denoising. Additionally, we further assess performance on other restoration tasks and evaluate zero-shot capabilities in the Sec. D.2.
@@
Dense image prediction tasks are evaluated across three vision tasks: ADE20k (2017b) for semantic segmentation, NYUv2 (2012) and SUNRGB-D (2014) for monocular depth estimation, and NYU-Depth v2 (2012) for surface normal estimation. Implementation details can be found in Sec. D.1.
Results. Table 1 presents a comprehensive performance comparison with recent state-of-the-art (SOTA) task-specific and all-in-one methods. As shown in the results, despite denoising and deraining data making up only a small portion of the overall training set, our method outperforms other unified methods and compare to some task-specific approaches. In the image grounding task, PixWizard significantly outperforms the diffusion-based generalist model InstructDiffusion by 4.8 cIoU on RefCOCO (val). However, there is still room for improvement compared to other highly specialized models. Furthermore, as shown in Fig.5, PixWizard supports flexible instructions, allowing it to highlight and visualize the target object directly on the image while also generating the corresponding binary mask. Additional quantitative evaluation results are provided in the Sec. D.3.
@@
For dense prediction tasks, PixWizard performs competitively against both generalist and task-specific baselines across all three tasks. In depth estimation on the NYUv2 test set, PixWizard achieves a 10.0% improvement in RMSE compared to Unified-IO and performs similarly to Painter and InstructCV. Additionally, Fig. 6 provides examples of PixWizard’s outputs. As shown, by supplying the corresponding task-specific prompt for the same image, we can easily generate the respective condition visualization, underscoring PixWizard’s significant practical value.
@@ 4.2 Second Section Results (Image Editing)
Settings. We evaluate PixWizard across two benchmarks, the MagicBrush Test (Zhang et al., 2024a) and the Emu Edit Test (Sheynin et al., 2024), to assess its effectiveness in image editing capabilities. For a fair comparison, we primarily compare it with several instruction-guided image editing methods. Consistent with Emu Edit, we use L1 distance, CLIP image similarity, DINO similarity, CLIP text-image similarity, and CLIP text-image direction similarity as metrics.
@@
Results. Table 2 presents our results compared to the baselines. The findings show that our model consistently outperforms InstructPix2Pix, MagicBrush, and UltraEdit in automatic metrics and achieves comparable performance to state-of-the-art method Emu Edit. As shown in Fig. 7, our model precisely identifies the editing region while preserving the rest of the pixels, and it demonstrates the best understanding of the given instructions.
@@ 4.3 Third Section Results (Image Generation)
In this section, we focus on evaluating the effectiveness of PixWizard’s generation capabilities. Specifically, we assess its performance across four tasks: text-to-image generation, controllable image generation, inpainting, and outpainting. Implementation details can be found in Sec. D.1.
@@
Controllable Generation Results. Without the need for separate training for each model, PixWizard is an all-in-one solution capable of handling multiple conditions. As shown in Table 3 and Fig. 8, PixWizard achieves the highest controllability and best image quality under depth conditions, while also being comparable to current separate models in image-text alignment.
Inpainting Results. As shown in Table 3, PixWizard outperforms other inpainting approaches, improving overall image quality based on FID and LPIPS scores. The qualitative examples in Fig. 9 further demonstrate PixWizard’s effectiveness in generating coherent content. Building on its strong inpainting capabilities, PixWizard enables users to perform more precise image editing tasks, such as Remove Anything, Replace Anything, and Add Anything. More details can be found in Sec. D.4
@@
Text-to-Image Results. As the results shown in Table 3, PixWizard achieves a FID score of 9.56 when tested for zero-shot performance on the COCO dataset. Although some models achieve a lower FID, they focus solely on text-to-image tasks and rely on significantly more training resources. Additionally, we evaluated the Human Preference Score (HPS v2), a robust benchmark for assessing human preferences in text-to-image synthesis. PixWizard performed well, delivering image quality comparable to popular text-to-image generators. We provide visual samples in Fig. 9. PixWizard supports high-resolution image synthesis, up to $1024\times 1024$, with any resolution and aspect ratio.
Figure 9: Visualization results of Inpainting, Outpainting and Text-to-Image Generation.